Morning Litho status report 2/28

Gary Sosa gsosa at stanford.edu
Mon Feb 28 07:30:45 PST 2011


Good morning All....

     Here is the morning Litho status report.

SVG Coat !- User reports that HPO transfer belts not turning. Will 
investigate on morning rounds

Karl Suss 1- User reports that system is not working. Will check out on 
morning rounds

Karl Suss 2- User reports that X/Y stage not moving well. Will check out 
on morning rounds

Microscope #2- Turret not turning with buttons. Will report to our 
microscope service people. Turret
                              can be rotated manually.

EV Bonder- Converting pressure plate from Steel to Quartz this morning.


     Please check back in Coral for latest tool updates and current 
equipment status.

        Thanks....  The Litho Team


*/Photolithography Area Status/*

	

	

*/Monday, February 28, 2011/*

	

	

	

	

*/General Information / Lab Activities/*

	

	

	

EV Bonder converting to Quartz pressure plate this morning

*/Exposure Tools/*

	

Status

	

Comments

	

Priority

ASML Stepper

	

*UP*

	

	

EVAlign

	

*PROBLEM*

	

Non-criticals only

	

4

EVAlign2

	

*UP*

	

	

EVG-Imprint

	

*UP*

	

	

Karl Suss 1

	

*PROBLEM*

	

System not responding

	

2

Karl Suss 2

	

*DOWN*

	

X/Y stage movement

	

1

*/Manual Coaters/*

	

Status

	

	

EVGSpraycoat

	

*UP*

	

	

Headway

	

*UP*

	

	

Laurel

	

*UP*

	

	

*/Coaters/*

	

Process Status

	

	

SVGCoat

	

3612 Process

	

*Problem*

	

Hot plate transport problem

	

3

220-3 Process

	

*Up*

	

	

220-7 Process

	

*Problem*

	

Optimizing Process

	

5

SVGCoat2

	

3612 Process

	

*Up*

	

	

955-0.7 Process

	

*Up*

	

	

3617 Process

	

*Up*

	

	

*/Developers/*

	

Status

	

	

SVGDEV

	

*UP*

	

	

SVGDEV2

	

*UP*

	

	

*/Ovens / Hotplates/*

	

Status

	

	

BLUEM

	

*UP*

	

	

Blue M 90 C.

	

*UP*

	

	

Blue M 110 C

	

*UP*

	

	

Blue M 150 C

	

*UP*

	

	

Fusion

	

*DOWN*

	

Hotplate experiments

	

UV-Cure

	

*UP*

	

	

White-Oven

	

*UP*

	

	

YES Vapor Prime

	

*UP*

	

	

*/Hotplates- /*

	

	

	

Laurel-left

	

*UP*

	

	

Laurel- right

	

*UP*

	

	

Headway

	

*UP*

	

	

Torrey Pines- left

	

*UP*

	

	

90 Deg C

	

*UP*

	

	

115 Deg C

	

*UP*

	

	

Torrey Pines- Right

	

*UP*

	

	

*/Metrology/Inspection/*

	

Status

	

	

Alpha-Step 500

	

*UP*

	

	

Nanospec 2-Litho

	

*UP*

	

	

Nanospec 1-Etch

	

*UP*

	

	

P2 Profiler

	

*PROBLEM*

	

-11 stage error

	

6

*/Microscpoes-/*

	

	

	

#1 Leica/Analytical

	

*UP*

	

	

#2 Ergolux/Drytek2

	

*PROBLEM*

	

Turret not rotating

	

7

#3 Olympus/Litho

	

*UP*

	

	

#4 Nikon Optiphot/Litho

	

*UP*

	

	

#5 Nikon/Litho

	

*UP*

	

	

#6 Leitz/Bosch

	

*UP*

	

	

#7 Leitz/Diffusion

	

*UP*

	

	

#8 Olympus/Litho

	

*UP*

	

	

#9 Olympus/Litho

	

*UP*

	

	

*/Miscellaneous/*

	

Status

	

	

EVBond

	

*UP*

	

	

KSBonder

	

*PROBLEM*

	

Leaking nitrogen hose

	

8

KSCoat

	

*UP*

	

	

Litho SRD

	

*UP*

	

	

Mask Cleaner

	

*UP*

	

	

	

	

	



-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/litho/attachments/20110228/b773a34f/attachment.html>


More information about the litho mailing list